Your first job out of college will pay you $73,000 in year 1 (exactly one year from today). You estimate that your salary will grow at 6% per year. You plan to retire in 37 years (you'll receive 37 years of salary). If the applicable discount rate is 9%, what is the present value of these future earnings today? Round to the nearest cent.
Answer #1

PV of annuity with growth = (P/ (r-g)) * (1- ((1+g)/(1+r))^n)

= (73000/(9%-6%))*(1-((1+6%)/(1+9%))^37

= 2433333.33* 0.643929=

1566893.93